OpenStudy (anonymous):

Find the exact value of the volume of a can of oil with a diameter 10 centimeters and height of 14 centimeters. Include correct units with your solution. Use "pi" in place of the pi symbol

OpenStudy (anonymous):

\[r=d/2=10/2=5\] \[volume= \pi*r^2*h\] \[=\pi*25*14\] \[=350\pi\]

OpenStudy (anonymous):

unit is\[cm^2\]

OpenStudy (asadkarim7):

cylinder area of base = pix25 height=14 vol of cylinder = area of bse xheight = 25pi x 14 350 pi

OpenStudy (anonymous):

pi*(d/2)^2 h = pi * 5^2 * 14 =350pi

OpenStudy (anonymous):

@mathematicalpro volume unit in cm^2 and not cm^3 ????????????

OpenStudy (anonymous):

ik and also written in the top